OpenStudy (anonymous):

I did a step by step algebraic solution in math , but could you write the algebraic solution in words because my english is bad A crate filled with oranges weighs 10 pounds The track team ate two-fifths of the oranges. The remaining oranges and crate together weighed 7 pounds. Find the weight of the empty crate.

OpenStudy (anonymous):

Please helpppp

OpenStudy (anonymous):

let the weight of crate = x and wt of oranges = y when 2/5 of oranges eaten, left is 3/5 of oranges = 3y/5 we are given x + y = 10 ------- (1) after 2/5 oranges eaten, then we are given x + 3y/5 = 7 ------- (2) subtract (2) from (1) and x is eliminated, giving y - 3y/5 = `10-7 (5y - 3y)/5 = 3 2y/5 = 3 2y = 3*5 y = 15/2 = 7.5 putting y = 7.5 in (1), we get x + 7.5 = 10 x = 10-7.5 x = 2.5 so wt of crate is 2.5 pounds

OpenStudy (anonymous):

if u r happy with the answer, pls click GOOD ANSWER button for me :))

OpenStudy (anonymous):

thank you veryyy much

OpenStudy (anonymous):

most and always welcome {{{ :)